### v1

# Robust Decision Aggregation with Adversarial Experts

Abstract
We consider a binary decision aggregation problem in the presence of both truthful and adversarial experts. The truthful experts will report their private signals truthfully with proper incentive, while the adversarial experts can report arbitrarily. The decision maker needs to design a robust aggregator to forecast the true state of the world based on the reports of experts. The decision maker does not know the specific information structure, which is a joint distribution of signals, states, and strategies of adversarial experts. We want to find the optimal aggregator minimizing regret under the worst information structure. The regret is defined by the difference in expected loss between the aggregator and a benchmark who makes the optimal decision given the joint distribution and reports of truthful experts.
We prove that when the truthful experts are symmetric and adversarial experts are not too numerous, the truncated mean is optimal, which means that we remove some lowest reports and highest reports and take averaging among the left reports. Moreover, for many settings, the optimal aggregators are in the family of piecewise linear functions. The regret is independent of the total number of experts but only depends on the ratio of adversaries. We evaluate our aggregators by numerical experiment in an ensemble learning task. We also obtain some negative results for the aggregation problem with adversarial experts under some more general information structures and experts’ report space.
1 Introduction
A company is preparing to launch a new product. To estimate demand, they assembled a focus group of five industry experts. The underlying true demand in the world can be in one of two states - high or low. Each expert reviewed available market data separately and made a binary prediction on whether demand will be high or low. For instance, three experts’ predictions are high (Expert 2, 4, 5) and two are low (Expert 1, 3). Suppose the experts are symmetric, i.e., with equal expertise. The company’s marketing team aggregated the predictions by taking simple averaging: since 3 out of 5 experts predicted high demand, they forecast a 60% chance of high demand. The accuracy of the aggregated probabilistic forecast can be evaluated using loss functions compared to the ground truth.
While simple averaging provides a practical approach to aggregating predictions from multiple experts, it does not always yield the optimal probabilistic forecast. With knowledge of the joint distribution between experts’ predictions and the true world state, more sophisticated Bayesian aggregation methods could potentially improve accuracy. The joint distribution is also called the information structure. However, in many cases the underlying information structure is unknown, and averaging has the advantage that it does not rely on explicit statistical assumptions about experts’ correlations. Prior research has shown averaging to be optimal under certain evaluation paradigms without knowledge of specific joint distributions (Neyman and Roughgarden,, 2022; Arieli et al.,, 2023).
In detail, they use the robust aggregation paradigm. There are two players playing a zero-sum game. Nature can choose an information structure $\theta$ from a family $\Theta$ and the decision maker needs to pick an aggregator $f$. Given an aggregator $f$ and the information structure $\theta$, nature’s payoff function is the regret $R(f,\theta)=\mathbb{E}_{\theta}[loss(f({\bm{x}}))-loss(opt({\bm{x}}))]$. It is the difference in the expected loss of $f$ compared to a benchmark aggregator $opt$. Specifically, $opt$ is an omniscient aggregator who knows the underlying information structure and outputs a forecast minimizing the expected loss. The decision maker aims to find the optimal aggregator to solve a minimax problem:
$\min_{f}\max_{\theta\in\Theta}R(f,\theta).$
Prior studies that employ the above paradigm typically assume all experts are truthful or rational. However, in reality, some experts may be influenced by external incentives that distort their predictions. For example, imagine all the experts make the correct forecast high, but competitors bribed two of the five experts in our scenario to deter the product launch. Despite seeing data indicating high demand, Experts 1 and 3 reported low demand due to bribery. Meanwhile, Experts 2, 4, and 5 provided uncorrupted high-demand predictions. The marketing team aggregated the predictions by taking simple averaging: since 3 out of 5 experts predicted high demand, they forecast a 60% chance of high demand. This estimate was biased from the true state by the manipulated opinions of 1 and 3. If the company had only listened to the uncorrupted opinions from 2, 4, 5, they would have correctly predicted 100% high demand. With the corrupted predictions, the aggregated forecast understated the true market demand. This example demonstrates a case where simple averaging can be misleading due to corrupted expert opinions. Such concerns also exist in other scenarios, e.g. when the jury debates on a verdict, when the judges rate the gymnasts, and when miners reach a consensus on a blockchain system. It raises an important research question:
Can we develop aggregation methods that are robust to corrupted experts, while maintaining simplicity and avoiding strong assumptions about information structures?
We find that the truncated mean is optimal under some mild assumptions for binary decision aggregation. If the marketing team knows there are at most two corrupted experts, they could remove two experts predicting low and two experts predicting high simultaneously and use simple averaging among the left experts. It will output 100% (1 out of 1) in our example. This method is also commonly used in the robustness fields, as well as some practical applications such as traditional sports.
Formally, to handle our question, it is natural to extend the robust aggregation paradigm by endowing nature the ability to choose not only the information structure but also the predictions of some experts. Suppose there are two types of experts, the truthful experts who always report truthfully and the adversarial experts who can choose a strategy $\sigma$ from a strategy space $\Sigma$. In this paper, we want to solve the minimax problem:
$\min_{f}\max_{\theta\in\Theta,\sigma\in\Sigma}R(f,\theta,\sigma)=\min_{f}\max_{\theta\in\Theta,\sigma\in\Sigma}\mathbb{E}_{\theta}[loss(f(\sigma({\bm{x}})))-loss(opt({\bm{x}}_{T}))].$
${\bm{x}}_{T}$ is the reports of truthful experts and $opt$ is the benchmark that knows the information structure $\theta$. Notice that the original robust forecast paradigm is a special case where the number of adversarial experts is zero. In this paper, we consider the adversarial experts who can report arbitrarily. But by setting the family of strategies $\Sigma$, we can also capture biased experts following traditional behavior models such as the Dawid-Skene model in the crowdsourcing field. Under this paradigm, we can generate different robust aggregators according to the features of applications.
1.1 Summary of Results
We aim to find the optimal aggregator in the presence of uncertain information structures and adversarial experts.
Binary Decision Aggregation
We first consider the binary decision aggregation problem (Arieli et al.,, 2023). In the original non-adversarial setting, each agent will see a binary private signal ($L$ or $H$) about the binary world state (0 or 1) and is asked to report their signal. We assume the experts are truthful and symmetric. That is, they share the same marginal distribution for signals and will report truthfully according to their private signals. The decision maker can only see anonymous reports and wants to forecast the probability of the world state 1.
In the adversarial setting, in addition to truthful experts, there exist some adversarial experts who will report arbitrarily from a choice set $\{H,L\}$. We endow the adversaries with the ability to observe the reports of others and collude.
We discuss two types of loss function, the L1 loss $\ell_{1}(x,y)=|x-y|$ and the L2 loss $\ell_{2}(x,y)=(x-y)^{2}$. Under L1 loss, the benchmark function $opt$ should report the maximum likelihood, while under L2 loss is the Bayesian posterior. Thus L1 loss is more suitable when we want to output a decision and L2 loss is used to output a probabilistic forecast. Suppose there are $n$ experts in total and $k=\gamma n$ adversarial experts. The theoretical results are shown in Table 1. The optimal aggregators are illustrated in Figure 1.
[…]
L1 Loss Non-adversarial $c$ Yes (Arieli et al.,, 2023)
Adversarial $c+O(\frac{\gamma}{1-2\gamma})$ Yes
L2 Loss Non-adversarial $c-O(\frac{1}{n})$ No, $O(\frac{1}{\epsilon})$$\star$
Adversarial $c+O(\frac{\gamma}{(1-2\gamma)^{2}})$ Yes
  • $\dagger$
    $c$ is a constant, $\gamma$ is the ratio of adversarial experts and $n$ is the number of experts.
  • $\star$
    The complexity for computing the $\epsilon$-optimal aggregator.
Table 1: Overview of main results.
0011112222333344445555666677778888999910101010000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111Number of Experts Reporting HOutput of AggregatorsOptimal AggregatorNon-adversarial: Optimal Aggregator, L1 lossAdversarial: Optimal Aggregator, L1 lossAdversarial: Optimal Aggregator, L2 lossNon-adversarial: Optimal Aggregator, L2 loss
Figure 1: Illustration of optimal aggregators for binary decision aggregation. All the aggregators are piecewise linear functions. For the adversarial setting, the optimal aggregators are both in the family of hard sigmoid functions.
  • L1 Loss Setting
    • Adversarial. We prove that the optimal aggregator is $k$-ignorance random dictator under the L1 loss, where $k$ is the number of adversarial experts. That is, we drop $k$ lowest reports as well as $k$ highest reports and randomly follow a left agent. It echoes the results in arieli2023universally that random dictator is the optimal aggregator in the non-adversarial setting. We also infer that the regret increases asymptotically linear with adversarial ratio $\gamma$ when $\gamma$ is small. Moreover, the regret is independent of the number of experts $n$. When $\gamma\approx 1/2$, the regret reaches the maximum $1/2$.
  • L2 Loss Setting
    • Non-adversarial. In this setting, the optimal aggregator is still piecewise linear, with the separation points in $\{1,n-1\}$. Although we do not obtain the closed-form for the optimal aggregator, we prove that we can compute the $\epsilon$-optimal aggregator in time $O(1/\epsilon)$. Unlike the L1 loss, the regret under the L2 loss is not independent of the number of experts $n$. It will increase with $n$. Compared to the L1 loss, the aggregator under L2 loss is conservative (i.e. closer to 1/2).
    • Adversarial. We provide a closed-form for the optimal aggregator which is also a hard sigmoid function with separation points in $\{k,n-k\}$ when $\gamma$ is small. The regret also increases asymptotically linear with parameter $\gamma$ and is independent of the number of experts $n$. When $\gamma\approx 1/2$, the regret reaches the maximum $1/4$. Interestingly, when we set $\gamma=0$, the formula of the adversarial optimal aggregator cannot induce the non-adversarial optimal aggregator. When $\gamma>0$, we can construct an equilibrium $(f,\theta,\sigma)$ such that the information structure $\theta$ has a zero loss benchmark. But when $\gamma=0$, we cannot construct an equilibrium $(f,\theta)$ in the same way as there is no adversarial strategy $\sigma$. The worst information structure in the equilibrium may have a non-zero loss benchmark, which leads to a more complex optimal aggregator.
Numerical Results
In Section 5, we evaluate our aggregators empirically in an ensemble learning task. In this task, we want to aggregate the output of some symmetric image classifiers. These classifiers are trained by different subsets of the full training set. We pick the cifar-10 datasets (Krizhevsky et al.,, 2009). Our experiment shows the optimality of our aggregator, and the majority vote is also a good approximation in practice. Under L2 loss, our aggregator has more advantages. The reason can be that the L1 loss only cares about the right answer while the L2 loss requires the confidence of the answer.
Extension to General Model
In Section 6, we extend the binary decision aggregation to the general model. In detail, we consider the general family of information structures and experts’ report space. For the general case, we provide negative results that a small group of adversaries can effectively attack the aggregator. Then we provide an important metric to estimate the regret. Intuitively, the metric is the maximal changes $k$ experts can make regarding the benchmark aggregator $opt$. The minimal regret can be bounded by this metric, with the help of a regularization parameter of information structures.
Main Technique
We solve the minimax problem by reducing the strategies spaces of adversaries and aggregators iteratively.
  • The adversarial side: We decompose the information structures into linear combinations of “simple” information structures. “Simple” means that the adversaries only use pure strategies and the report space is small. Due to the convexity of the regret function, the maximum value is always given by these “simple” information structures.
  • The decision maker’s side: For each “simple” information structure, it is easy to find the best aggregator responding to it. We pick the aggregator with the maximum regret among the best response aggregators and extend it to a full aggregator by linear interpolation. Then we prove that this extended function is optimal.
2 Related Work
Robust Information Aggregation
Arieli et al., (2018) first propose a robust paradigm for the information aggregation problem, which aims to minimize the regret under the worst information structures. They mainly study the conditional independent setting and provide the average prior aggregator. There is a growing number of research on the robust information aggregation problem. Neyman and Roughgarden, (2022) also use the robust regret paradigm under the projective substitutes condition and shows that choosing a random expert is asymptotically optimal. De Oliveira et al., (2021) consider the robust absolute paradigm and prove that we should pay more attention to the best single information source. Pan et al., (2023) consider the optimal aggregator with second-order information. Guo et al., (2024) provide an algorithmic method to compute the near-optimal aggregator for the conditional independent information structure. In this paper, we mainly consider a different information structure and obtain an exact optimal aggregator.
Our paper is most relevant to Arieli et al., (2023), which considers the symmetric agents setting with the same marginal distribution. They prove the random dictator strategy is optimal under some mild conditions. We extend their results to different loss functions and the adversarial setting. Our results show that the optimal aggregator is piece-wise linear, sharing similar properties to the random dictator under several different circumstances.
Adversarial Information Aggregation
In the crowdsourcing field, some works aim to detect unreliable workers based on observed labels. They mainly consider two kinds of unreliable workers, the truthful workers but with a high error rate, and the adversarial workers who will arbitrarily assign labels. One possible approach is using the “golden standard” tasks, which means managers know the ground truth (Snow et al.,, 2008; Downs et al.,, 2010; Le et al.,, 2010). When there are no “golden standard” tasks, Vuurens et al., (2011); Hovy et al., (2013); Jagabathula et al., (2017); Kleindessner and Awasthi, (2018) detect the unreliable workers via revealed labels under some behavior models such as the Dawid-Skene model. Other works focus on finding the true labels with adversarial workers. Steinhardt et al., (2016) consider a rating task with $\alpha n$ reliable workers, and others are adversarial workers. They showed that the managers can use a small amount of tasks, which is not scaled with $n$, to determine almost all the high-quality items. Ma and Olshevsky, (2020) solve the adversarial crowdsourcing problem by rank-1 matrix completion with corrupted revealed entries. Other works also focus on the data poison attacks in crowdsourcing platforms (Chen et al.,, 2021, 2023; Fang et al.,, 2021; Gadiraju et al.,, 2015; Miao et al.,, 2018; Yuan et al.,, 2017). Unlike crowdsourcing, we do not assume the existence of many similar tasks and focus on the one-shot information aggregation task.
Han et al., (2021) and Schoenebeck et al., (2021) propose a peer prediction mechanism for a hybrid crowd containing both self-interested agents and adversarial agents. Han et al., (2021) focus on the binary label setting and propose two strictly posterior truthful mechanisms. Schoenebeck et al., (2021) prove their mechanism guarantees truth-telling is a $\epsilon$-Bayesian Nash equilibrium for self-interested agents. The focus of our paper is the aggregation step, thus we do not consider the incentives of non-adversarial agents but assume they are truth-telling.
Kim and Fey, (2007) study the voting problem when there are voters with adversarial preferences. They prove that it is possible that a minority-preferred candidate almost surely wins the election under equilibrium strategies. They want to determine when the majority vote can reveal the true groundtruth while we want to find a robust aggregator for any situation.
Robust Ensemble Learning
Ensemble learning methods leverage the power of multiple models to improve prediction accuracy, similar to how aggregating predictions from a diverse crowd can produce better forecasts than individual opinions. The earliest work of ensemble learning can date back to the last century (Dasarathy and Sheela,, 1979). They aim to combine different classifiers trained from different categories into a composite classifier. schapire1990strength provides a new algorithm to convert some weak learners to strong learners. The most widely-used ensemble learning methods include bagging (breiman1996bagging), AdaBoost (hastie2009multi), random forest (breiman2001random), random subspace (ho1995random), gradient boosting (friedman2002stochastic). dong2020survey provides a comprehensive review of ensemble learning. The main difference between ensemble learning and information aggregation is that ensemble learning is a training process and thus involves multi-round aggregation. Unlike them, our method does not assume any knowledge about the underlying learning models but only needs the final outputs of models.
In the robust learning field, there are many works aimed at data poison attacks to improve the robustness of learning algorithms (jagielski2018manipulating; zhao2017efficient; biggio2012poisoning). In comparison, our adversaries cannot change the training data, but alter the output of learning models.
3 Problem Statement
{toappendix}
We use denote $\{0,1,\cdots,n\}$ as $[n]$ and $\Delta_{X}$ as all possible distributions of $X$. For a distribution $P$, we use $supp(P)$ to represent its support.
Suppose the decision maker (DM) wants to determine the true state $\omega$ from a binary choice set $\Omega=\{0,1\}$, e.g. is the demand low or high? She is uninformed and asks $n$ experts for advice. Each expert will receive a signal $s_{i}$ from a binary space $\{L,H\}$ (low, high), indicating that with low (high) probability the state is 1. Experts share the same prior $\mu=\Pr[\omega=1]$ and posterior $p_{0}=\Pr[\omega=1|s_{i}=L]$ and $p_{1}=\Pr[\omega=1|s_{i}=H]$. Without loss of generality, we assume $p_{0}<\frac{1}{2}<p_{1}$. Let $a=\Pr[s_{i}=H|\omega=1],b=\Pr[s_{i}=H|\omega=0]$, which is the conditional probability of receiving signal H. The joint distribution $\theta$ of $\omega,s_{i}$ is drawn from a family $\Theta$.
In the non-adversarial setting, we assume the experts are truthful, which means they will truthfully report their signals. We could relax this assumption to rational experts who aim to maximize their revenues by proper incentive mechanisms after the true state is revealed. Since the experts are symmetric to the decision maker and the choice is binary, the reports of experts can be compressed by the number of experts reporting $H$, denoted by $X=\sum_{i=1}^{n}\mathbbm{1}(X_{i}=H)$. The output of DM can be represented by a real number in the range $[0,1]$, indicating her belief about state 1. Thus DM needs to choose an aggregator $f:[n]\to[0,1]$.
In the adversarial setting, there are $k=\gamma n$ adversarial experts who can arbitrarily report from $\{L,H\}$. We assume they are omniscient. That is, they know the true state and reports of other truthful experts and can collude. Their strategy is represented by $\sigma\in\Sigma$, which is a mapping from $\{0,1\}\times[n-k]$ to a distribution in $\Delta_{[k]}$. Suppose the set of truthful experts is $T$, and the set of adversarial experts is $A$. We denote the number of truthful experts reporting $H$ by $X_{T}=\sum_{i\in T}\mathbbm{1}(X_{i}=H)$. Similarly, we have $X_{A}=\sum_{i\in A}\mathbbm{1}(X_{i}=H)$ which is the number of adversarial experts reporting $H$.
Regret Robust Paradigm
Given a family of joint distributions $\Theta$ and a family of strategies $\Sigma$, DM aims to minimize the regret compared to the non-adversarial setting in the worst information structure. That is, DM wants to find the optimal function $f^{*}$ to solve the following min-max problem:
$R(\Theta,\Sigma)=\inf_{f}\sup_{\theta\in\theta,\sigma\in\Sigma}\mathbb{E}_{\theta,\sigma}[\ell(f(x),\omega)]-\mathbb{E}_{\theta}[\ell(opt_{\theta}(x_{T}),\omega)].$
where $opt_{\theta}(\cdot)$ is a benchmark function, outputting the optimal result given the joint distribution $\theta$ and truthful experts’ reports $x_{T}$ to minimize the expected loss:
$opt_{\theta}(x_{T})=\operatorname*{arg\,min}_{g}\mathbb{E}_{\theta}[\ell(g(x_{T}),\omega)].$
$\ell$ is a loss function regarding the output of the aggregator and the true state $\omega$. We consider two commonly used loss function, L1 loss (or absolute loss) $\ell_{1}(y,\omega)=|y-\omega|$ and L2 loss (or square loss) $\ell_{2}(y,\omega)=(y-\omega)^{2}$. L2 loss will punish the aggregator less when the prediction is closer to the true state. Thus it encourages a more conservative strategy for the aggregator to improve the accuracy in the worst case. L1 loss will encourage a radical strategy to approximate the most possible state.
For short, we also define $R(f,\theta,\sigma)=\mathbb{E}_{\theta,\sigma}[\ell(f(x),\omega)]-\mathbb{E}_{\theta}[\ell(opt_{\theta}(x_{T}),\omega)]$ is aggregator $f$’s regret on information structure $\theta$ and adversarial strategy $\sigma$. $R(f,\Theta,\Sigma)=\sup_{\theta\in\Theta,\sigma\in\Sigma}R(f,\theta,\sigma)$ is the maximal regret of aggregator $f$ among the family $\Theta,\Sigma$.
3.1 Auxiliary Tools
Now we define two important distributions $v_{1},v_{0}\in\Delta_{[n]}$, which are the distributions of $X$ conditioning on the world state: $v_{1}(x)=\Pr_{\theta,\sigma}[X=x|\omega=1],v_{0}(x)=\Pr_{\theta,\sigma}[X=x|\omega=0]$. Notice that $v_{1},v_{0}$ are related to both $\theta$ and $\sigma$. Similarly, we define $u_{1}(x)=\Pr_{\theta}[X_{T}=x|\omega=1],u_{0}(x)=\Pr_{\theta}[X_{T}=x|\omega=0]$ which are the conditional distributions of truthful experts. Notice that $u_{1},u_{0}\in\Delta_{[n-k]}$.
The expected loss of the aggregator can be decomposed by $v_{1},v_{0}$:
$\mathbb{E}_{\theta,\sigma}[\ell(f(x),\omega)]=\mu E_{x\sim v_{1}}[\ell(f(x))]+(1-\mu)E_{x\sim v_{0}}[\ell(1-f(x))].$
By simple calculation, we can obtain the closed-form of the benchmark $opt_{\theta}(x)$ under different loss function $\ell$.
L1 loss
$opt_{\theta}(x_{T})=\left\{\begin{aligned} 1&\quad\mu u_{1}(x_{T})\geq(1-\mu)u_{0}(x_{T})\\ 0&\quad otherwise\\ \end{aligned}\right.$
The expected loss is $1-\sum_{x}\max(\mu u_{1}(x),(1-\mu)u_{0}(x))$.
L2 loss
$opt_{\theta}(x_{T})=\frac{\mu u_{1}(x_{T})}{\mu u_{1}(x_{T})+(1-\mu)u_{0}(x_{T})}.$
The expected loss is $\sum_{x}\frac{\mu(1-\mu)u_{1}(x)u_{0}(x)}{\mu u_{1}(x)+(1-\mu)u_{0}(x)}$
Before considering the optimal aggregator, we first characterize all feasible information structures. Section 3.1 fully formalizes the distribution $u_{1},u_{0}$, i.e. the possible reports of $n-k$ truthful experts.
{lemma}
[(arieli2023universally)] For any distribution $u_{1},u_{0}\in\Delta_{[n-k]}^{2}$, there exists an information structure $\theta$ such that $u_{1}(x)=\Pr_{\theta}[X_{T}=x|\omega=1],u_{0}(x)=\Pr_{\theta}[X_{T}=x|\omega=0]$ if an only if $\mathbb{E}_{x\sim u_{1}}[x]=(n-k)a$ and $\mathbb{E}_{x\sim u_{0}}[x]=(n-k)b$.
The following lemma is an extension of section 3.1. It shows that there is also some restriction on the expectation of the distribution of inputs for adversarial experts and truthful experts. The proof is deferred to Section 3.
{lemmarep}
For any information structure $\theta$ and adversarial strategy $\sigma)$, the corresponding $v_{1},v_{0}$ satisfy $(n-k)a\leq\mathbb{E}_{x\sim v_{1}}[x]\leq(n-k)a+k$ and $(n-k)b\leq\mathbb{E}_{x\sim v_{0}}[x]\leq(n-k)b+k$.
{appendixproof}
By 3.1, $\mathbb{E}_{x\sim u_{1}}[x]=(n-k)a$ and $\mathbb{E}_{x\sim u_{0}}[x]=(n-k)b$. Consider $\mathbb{E}_{x\sim v_{1}}[x]-\mathbb{E}_{x\sim u_{1}}[x]$, which is the expectation of reports of adversaries. Since there are at most $k$ adversaries. Then $0\leq\mathbb{E}_{x\sim v_{1}}[x]-\mathbb{E}_{x\sim u_{1}}[x]\leq k$. Thus $(n-k)a\leq\mathbb{E}_{x\sim v_{1}}[x]\leq(n-k)a+k$. Similarly, $(n-k)b\leq\mathbb{E}_{x\sim v_{0}}[x]\leq(n-k)b+k$.
4 Theoretical Results
In this section, we discuss the optimal aggregator under different settings.
4.1 L1 loss
We start from the optimal aggregator under L1 loss. In the non-adversarial setting, arieli2023universally prove the optimal aggregator is the random dictator, i.e. randomly and uniformly following an expert. It is equivalent to the simple averaging $f(x)=x/n$.
In the adversarial setting, intuitively, we need to be more radical in the corner case such as $x\leq k$ or $x\geq n$ since it is easy to be distorted and lacks confidence. Thus we will ignore the influence of the minor reports in these cases. In formal, we prove that the optimal aggregator is a hard sigmoid function from 0 to 1, which is called the $k$-ignorance random dictator when the ratio of adversaries is small. It can be explained by such a strategy that removes $k$ lowest reports and $k$ highest reports and then randomly and uniformly follows a left expert. Notice that $k$ is the number of adversarial experts.
Definition 4.1 ($k$-ignorance random dictator).
We call $f$ is $k$-ignorance random dictator if
$f(x)=\left\{\begin{aligned} &1&\quad x\geq n-k\\ &0&\quad x\leq k\\ &\frac{x-k}{n-2k}&\quad otherwise\\ \end{aligned}\right.$ (1)
Theorem 4.2.
When $\gamma\leq\min\left(\frac{a\mu-(1-\mu)b}{\mu+a\mu-(1-\mu)b},\frac{a}{1+a},\frac{1-b}{2-b}\right)$, the $k$-ignorance random dictator is optimal under the L1 loss. Moreover, the regret
$R(\Theta,\Sigma)=\mu+\frac{(1-\mu)((1-\gamma)b+\gamma)-\mu(1-\gamma)a}{1-2\gamma}.$
Proof Sketch
{toappendix}
Before we prove the main theorem, we need some lemmas first. Section 4.1 shows that there exists one worst information structure with reports in $[k,n-k]$. Section 4.1 shows in most cases, the optimal aggregator should be linear in $[k,n-k]$.
{lemma}
Let $(\Theta^{\prime},\Sigma^{\prime})=\{\theta,\sigma|\theta\in\Theta,\sigma\in\Sigma,k\leq X\leq n-k\}$. Then $R(\Theta,\Sigma)=R(\Theta^{\prime},\Sigma^{\prime})$.
Proof.
Since $R(\Theta,\Sigma)\geq R(\Theta^{\prime},\Sigma^{\prime})$, we only need to prove $R(\Theta,\Sigma)\leq R(\Theta^{\prime},\Sigma^{\prime})$.
For any $(\theta,\sigma)$ where exists $s^{\prime}$ such that $\sigma(s_{T}^{\prime})+x_{T}<k$, we find another $\sigma^{\prime}$ with $\sigma^{\prime}(s_{T}^{\prime})+x_{T}=k$. Notice that it is always possible since there are $k$ adversaries. We want to prove
$\min_{f}\max_{\theta,\sigma}R(f,\theta,\sigma)\leq\min_{f}\max_{\theta,\sigma^{\prime}}R(f,\theta,\sigma^{\prime}).$
Suppose $\operatorname*{arg\,min}_{f}\max_{\theta}R(f,\theta,\sigma^{\prime})=f^{*}$, let $f^{\prime}(\sigma(s_{T}^{\prime})+s_{T}^{\prime})=f^{*}(k)$ and other values are the same as $f^{*}$, then
$\max_{\theta,\sigma}R(f^{\prime},\theta,\sigma)=\max_{\theta,\sigma^{\prime}}R(f^{*},\theta,\sigma^{\prime})=\min_{f}\max_{\theta,\sigma^{\prime}}R(f^{*},\theta,\sigma^{\prime}).$
Thus $\min_{f}\max_{\theta}R(f,\theta,\sigma)\leq\min_{f}\max_{\theta}R(f,\theta,\sigma^{\prime})$. In finite steps, we can replace every $s^{\prime}$ in $\sigma$ with $\sigma(s_{T}^{\prime})+s_{T}^{\prime}<k$. So there exists $\sigma^{\prime}\in\Sigma^{\prime}$ such that $\min_{f}\max_{\theta}R(f,\theta,\sigma)\leq\min_{f}\max_{\theta}R(f,\theta,\sigma^{\prime})$. So $R(\Theta,\Sigma)\leq R(\Theta,\Sigma^{\prime})$.
By similar argument, we can let $f^{\prime}(\sigma(s_{T}^{\prime})+s_{T}^{\prime})=f^{*}(n-k)$ for those $\sigma(s_{T}^{\prime})+s_{T}^{\prime}>n-k$. So $R(\Theta,\Sigma)\leq R(\Theta,\Sigma^{\prime})$. ∎
{lemma}
Fix $f$ and let $l$ be the line passing through $(k,f(k))$ and $(n-k,f(n-k))$. Suppose there exist $x_{1},x_{2},x_{3},x_{4}\in\{k,k+1,\cdots,n-k\}$ s.t. $x_{1}\leq(n-k)a\leq x_{2},x_{3}\leq k+(n-k)b\leq x_{4}$ with $(x_{1},f(x_{1})),(x_{2},f(x_{2}))$ not above $l$ and $(x_{3},f(x_{3})),(x_{4},f(x_{4}))$ not below $l$. In addition, $x_{1},x_{2},x_{3}-k,x_{4}-k$ are distinct. Then
$R(f)\geq R(l)=\mu+(1-2\mu)f(k)+\frac{((1-\mu)(nb+k)-\mu na)(f(n-k)-f(k))}{n-2k}.$
Proof.
Suppose there are such $x_{1},x_{2},x_{3},x_{4}$, we select the information structure with support $supp(v_{1})=\{x_{1},x_{2}\},supp(v_{0})=\{x_{3},x_{4}\}$. It is possible since we can first select $\theta$ with support $supp(u_{1})=\{x_{1},x_{2}\},supp(u_{0})=\{x_{3}-k,x_{4}-k\}$ by section 3.1. The strategy of adversaries is to always report L when the state is 1, and report H when the state is 0. Since $x_{1},x_{2},x_{3}-k,x_{4}-k$ are distinct, the benchmark loss is 0. Now we consider $R(f,\theta,\sigma)$,
$R(f,\theta,\sigma)$ $=\mu(1-\mathbb{E}_{x\sim v_{1}}[f(x)])+(1-\mu)(\mathbb{E}_{x\sim v_{0}}[f(x)])$
$\geq\mu(1-\mathbb{E}_{x\sim v_{1}}[(n-2k-x)f(k)/(n-2k)+(x-k)f(n-k)/(n-2k)])$
$+(1-\mu)(\mathbb{E}_{x\sim v_{0}}[(n-2k-x)f(k)/(n-2k)+(x-k)f(n-k)/(n-2k)])$
The inequality holds because $(x_{1},f(x_{1})),(x_{2},f(x_{2}))$ not above $l$ and $(x_{3},f(x_{3})),(x_{4},f(x_{4}))$ not below $l$.
On the other hand, since $l$ is linear, we have
$R(l,\theta,\sigma)$ $=\mu(1-\mathbb{E}_{x\sim v_{1}}[l(x)])+(1-\mu)(\mathbb{E}_{x\sim v_{0}}[l(x)])$
$=\mu(1-\mathbb{E}_{x\sim v_{1}}[(n-2k-x)f(k)/(n-2k)+(x-k)f(n-k)/(n-2k)])$
$+(1-\mu)(\mathbb{E}_{x\sim v_{0}}[(n-2k-x)f(k)/(n-2k)+(x-k)f(n-k)/(n-2k)])$
By section 3.1,
$\mu(1-\mathbb{E}_{x\sim v_{1}}[(n-2k-x)f(k)/(n-2k)+(x-k)f(n-k)/(n-2k)])$
$=$ $\mu(1-f(k)+\frac{kf(n-k)}{n-2k}-\mathbb{E}_{x\sim v_{1}}[x]\frac{f(n-k)-f(k)}{n-2k})$
$\leq$ $\mu(1-f(k)+\frac{kf(n-k)}{n-2k}-(n-k)a\frac{f(n-k)-f(k)}{n-2k})$
and
$(1-\mu)(\mathbb{E}_{x\sim v_{0}}[(n-2k-x)f(k)/(n-2k)+(x-k)f(n-k)/(n-2k)])$
$=$ $(1-\mu)(f(k)-\frac{kf(n-k)}{n-2k}+\mathbb{E}_{x\sim v_{0}}[x]\frac{f(n-k)-f(k)}{n-2k})$
$\leq$ $(1-\mu)(f(k)-\frac{kf(n-k)}{n-2k}+((n-k)b+k)\frac{f(n-k)-f(k)}{n-2k})$
Thus,
$R(f)\geq R(l)=\mu+(1-2\mu)f(k)+\frac{((1-\mu)((n-k)b+k)-\mu(n-k)a)(f(n-k)-f(k))}{n-2k}.$
Now we start to prove our main theorem.
Proof of Theorem 4.2.
First by Section 4.1 we only need to consider the aggregators lying in $[k,n-k]$. We can pick $x_{1}=x_{3}=k,x_{2}=x_{4}=n-k$, by section 4.1 we know that the optimal aggregator should be linear in $[k,n-k]$. Notice that $R(l)$ is decreasing with $f(n-k)$ and increasing with $f(k)$, the minimum is given by $f(n-k)=1,f(k)=0$. Thus the optimal aggregator should be the $k$-ignorance dictator. And plug the value of $f(k),f(n-k)$ we obtain that
$R(\Theta,\Sigma)=\mu+\frac{((1-\mu)((n-k)b+k)-\mu(n-k)a)}{n-2k}.$
We provide a proof sketch here and the complete proof is deferred to Section 4.1. We find the optimal aggregator in two steps. First, we show there exists one worst information structure with reports always lying in $[k,n-k]$ for any $k<n/2$. Moreover, we characterize that the optimal aggregator is a constant when $x\geq n-k$ and $x\leq k$.
Then we prove that in $[k,n-k]$, there exists a linear optimal aggregator or we can find a linear aggregator with non-increasing regret. Thus the optimal aggregator is fully characterized and the only thing left is to find the optimal value of $f(k)$ and $f(n-k)$. By calculation, the minimum is given by 0 and 1 when $\gamma$ is small, which completes our proof.
Discussion
Notice that when $k=0$, the 00-ignorance random dictator is the same as the random dictator. Thus the non-adversarial setting is a special case of the adversarial setting where $k=0$. Intuitively, the $k$-ignorance strategy protects DM from adversaries with extreme reports, which is the optimal strategy for adversaries when the aggregator is monotone. It is also called the truncated mean and is commonly used to increase the robustness of the systems. In particular, it is often seen in many traditional sports such as gymnastics contests.
When $\gamma>\min\left(\frac{a\mu-(1-\mu)b}{\mu+a\mu-(1-\mu)b},\frac{a}{1+a},\frac{1-b}{2-b}\right)$, the $k$-ignorance random dictator may not be optimal. However, by the same argument, we can still infer that the optimal aggregator is piecewise linear. In fact, the optimal aggregator could be a constant when $\gamma$ is large, with a possible value of 1 or 0. It means that the aggregator is uninformative given any reports from the experts for large $\gamma$.
If the Number of Adversaries $k$ is Unknown
To determine the optimal aggregator, we need to know the parameter $k$. In practice, we may be unable to know the exact number of adversaries. Instead, we can only estimate the upper bound $K$ of $k$. In this case, we prove that the optimal aggregator is the $K$-ignorance random dictator (Section 4.1). The intuition is that when we increase the number of adversaries, the newly added adversaries can always pretend to be truthful experts such that the regret will not decrease. Thus most adversaries are always the worst case to handle. This result can also be extended to the L2 loss setting.
{lemmarep}
Suppose $(\Theta_{k},\Sigma_{k})$ includes all information structures and strategies with $n-k$ truthful experts and $k$ adversarial experts. Let $(\Theta,\Sigma)=\bigcup_{k\leq K}(\Theta_{k},\Sigma_{k})$, which includes all information structures and adversarial strategies with adversaries less than $K$. When $\frac{K}{n}\leq\min\left(\frac{a\mu-(1-\mu)b}{\mu+a\mu-(1-\mu)b},\frac{a}{1+a},\frac{1-b}{2-b}\right)$, the $K$-ignorance random dictator is optimal.
{appendixproof}
Fix $N$, for any $f$, let $R(k)=\mu+\frac{((1-\mu)(nb+k)-\mu na)}{n-2k}$. Then $R(f,\Theta_{k},\Sigma_{k})\geq R(k)$. Since $R(k)$ is increasing in $k$, we obtain that $R(k)=R(K)$. Thus for any $f$,
$R(f,\Theta,\Sigma)\geq\max_{k}R(f,\Theta_{k},\Sigma_{k})=R(K).$
On the other hand, For the $K$-ignorance random dictator $f^{K}$, $R(f,\Theta,\Sigma)=R(K)$, which shows it is optimal.
4.2 L2 loss
Now we consider the L2 loss $\ell_{2}$.
4.2.1 Warm-Up
We start from a warm-up case. In the L1 loss setting, the optimal aggregator is independent with three parameters we defined before, the prior $\mu$ and the expected posteriors $a,b$. It means the optimal aggregator is universal. However, in the L2 loss setting, the optimal aggregator will also depend on these parameters. When these parameters are unknown, it is impossible to obtain any informative aggregator (Section 4.2.1, Section 4.2.1).
{lemmarep}
When $\mu,a,b$ is unknown, the optimal aggregator is the random guess, that is, $f^{*}(x)=1/2$. Formally, let $(\Theta_{\mu,a,b},\Sigma_{\mu,a,b})$ includes all information structures and adversarial strategies with parameters $\mu,a,b$ and $(\Theta,\Sigma)=\bigcup_{0\leq\mu\leq 1,a>b}(\Theta_{\mu,a,b},\Sigma_{\mu,a,b})$. $R(\Theta,\Sigma)=1/4$. It holds for both adversarial and non-adversarial settings.
{appendixproof}
We first consider the non-adversarial setting. On the one hand, since $\ell(1/2)=1/4$, for any $\theta,\sigma$,
$R(f^{*},\theta,\sigma)$ $\geq\Pr_{\theta}[\omega=1]\ell(1/2)+\Pr_{\theta}[\omega=0]\ell(1/2)$
$=1/4$
Thus $R(f^{*},\Theta,\Sigma)=1/4$.
On the other hand, we select $x_{1}=0,x_{2}=1,x_{3}=n-1,x_{4}=n$, and $\mu=0.5$. Consider the mixture of two information structures with uniform distribution. In the first one, $supp(v_{1})=\{x_{1},x_{4}\},supp(v_{0})=\{x_{2},x_{3}\}$. In the second one, $supp(v_{1})=\{x_{2},x_{3}\},supp(v_{0})=\{x_{1},x_{4}\}$. Let $b=2/n$, consider a sequence $\{a_{t}\}$ such that $\lim_{t\rightarrow\infty}a_{t}=b$ and $a_{t}>b$.
By this sequence of parameter $a$ we construct a sequence of mixed information structures $\theta_{t}$. We will find that $\lim_{t\rightarrow\infty}R(\theta_{t})=1/4$ since the Bayesian posterior of $\theta_{t}$ is $f(x_{1})=f(x_{1})=f(x_{3})=f(x_{4})=1/2$. Thus $R(f,\Theta,\Sigma)\geq 1/4$ for any $f$.
For the adversarial setting, the random guess will still obtain regret $1/4$. the regret of other aggregators will not decrease. Thus we complete our proof for both adversarial and non-adversarial setting.
{lemmarep}
When $\mu$ is known but $a,b$ is unknown, then the optimal aggregator is the prior guess, that is $f^{*}(x)=\mu$. Formally, let $(\Theta_{a,b},\Sigma_{a,b})$ includes all information structures and adversarial strategies with parameters $\mu,a,b$ and $(\theta,\sigma)=\bigcup_{a>b}(\Theta_{a,b},\Sigma_{a,b})$. Then $R(\Theta,\Sigma)=\mu(1-\mu)$. It holds for both adversarial and non-adversarial settings.
{appendixproof}
Similarly, we first consider the non-adversarial setting. On the one hand,
$R(f^{*},\theta,\sigma)$ $\geq\Pr_{\theta}[\omega=1]\ell(\mu)+\Pr_{\theta}[\omega=0]\ell(1-\mu)$
$=\mu(1-\mu)$
Again, we select $x_{1}=0,x_{2}=1,x_{3}=n-1,x_{4}=n$. Consider the mixture of two information structures. In the first one, $supp(v_{1})=\{x_{1},x_{4}\},supp(v_{0})=\{x_{2},x_{3}\}$. In the second one, $supp(v_{1})=\{x_{2},x_{3}\},supp(v_{0})=\{x_{1},x_{4}\}$. Using the same argument in Section 4.2.1 we will obtain that for any $f$, $R(f,\theta,\sigma)\geq\mu(1-\mu)$. Thus $f^{*}$ is the optimal aggregator.
The non-trivial setting appears when all the parameters are known to the aggregator. In the rest of this section, we assume $\mu,a,b$ are known.
4.2.2 Non-adversarial setting
First, we discuss the non-adversarial setting. Unfortunately, the optimal aggregator is not hard sigmoid in L2 loss when there are no adversaries. Furthermore, we cannot obtain a closed-form of the optimal aggregator unless we can solve a polynomial equation of high order. However, we can still calculate it efficiently. We state the result as theorem 4.3.
Theorem 4.3.
There exists an algorithm that costs $O(1/\epsilon)$ to find the $\epsilon$-optimal aggregator in the non-adversarial setting. The $\epsilon$-optimal aggregator $f^{\prime}$ means that $R(f^{\prime},\Theta)\leq\min_{f}R(f,\Theta)+\epsilon$.
Proof Sketch
We prove this theorem in several steps. First, similar to the flow in arieli2018robust, we do the basic dimension reduction on $\Theta$, which allows us to consider information structures with at most 4 possible reports (Section 4.2.2).
{lemmarep}
Let $\Theta_{4}=\{\theta\in\Theta|supp(u_{1})\leq 2,supp(u_{0})\leq 2\}$. Then for any $f$, we have $R(f,\Theta,\Sigma)=R(f,\Theta_{4},\Sigma)$.
{appendixproof}
Consider the distribution vector $u_{1}$, it should satisfy the following two linear constraints:
$\sum_{x}u_{1}(x)=1$
$\sum_{x}xu_{1}(x)=na$
By the basic theorem of linear programming, we can decompose $u_{1}$ with a linear combination of some extreme distribution vectors. That is, there exists $t_{1},\cdots,t_{m}$ and $\lambda_{1},\cdots,\lambda_{m}$ such that
$u_{1}(x)=\sum_{i}\lambda_{i}t_{i}(x)$
$\sum_{x}t_{i}(x)=1$
$\sum_{x}xt_{i}(x)=na$
The extreme point here means that $supp(t_{i})\leq 2$ for any $i$. Now fix $u_{0}$, we can create a new information structure $\theta_{i}$ by setting the $t_{i}(x)=\Pr_{\theta_{i}}[X=x|\omega=1]$ and $u_{0}(x)=\Pr_{\theta_{i}}[X=x|\omega=0]$.
Since the benchmark aggregator $\frac{u_{1}(x)u_{0}(x)}{u_{1}(x)+u_{0}(x)}$ is concave with $u_{1}(x)$ given $u_{0}(x)$, and $\mathbb{E}_{\theta}[\ell(f)]$ is linear in $u_{1}$ given $f$. Thus for any $f$,
$R(f,\theta)\leq\sum_{i}\lambda_{i}R(f,\theta_{i})\leq\max_{i}R(f,\theta_{i}).$
Then following the same argument in $u_{0}$ we obtain that we can decompose $u_{0}$ with some extreme vectors with support space less than 2, which completes our proof for $R(f,\Theta)=R(f,\Theta_{4})$.
However, we still have around $\binom{n}{4}$ possible information structures. Furthermore, we show that we can only consider those “extreme information structures” with extreme report support space (Section 4.2.2). “Extreme” here means reports in $\{0,1,n-1,n\}$. Thus the meaningful information structures are reduced to constant size.
{lemmarep}
Let $\Theta_{e}=\{\theta\in\Theta_{4}|supp(u_{1})\subset\{0,1,n-1,n\},supp(u_{0})\subset\{0,1,n-1,n\}\}$. We have $R(\Theta,\Sigma)=R(\Theta_{e},\Sigma)$.
{appendixproof}
Consider the optimal aggregator $f^{*}=\arg\min_{f}R(f,\Theta_{e})$. We extend $f^{*}$ by linear interpolation in the non-extreme input:
$f(x)=\left\{\begin{aligned} &f^{*}(x)&\quad x=\{0,1,n-1,n\}\\ &\frac{f^{*}(n-1)-f^{*}(1)}{n-2}*(x-1)&\quad otherwise\\ \end{aligned}\right.$ (2)
Then for any $22\leq x\leq n-$ 2, we can write $x$ as linear combination of $1$ and $n-1$, $x=\alpha(x)*1+(1-\alpha(x))*(n-1)$ where $\alpha(x)=\frac{n-1-x}{n-2}\in[0,1]$. Since both $(1-f(x))^{2}$ and $f(x)^{2}$ is convex in $[1,n-1]$,
$\mathbb{E}_{\theta}[\ell(f)]$ $=\mu\sum_{x}u_{1}(x)(1-f(x))^{2}+(1-\mu)\sum_{x}u_{0}(x)f(x)^{2}$
$\leq\mu\sum_{x\in\{0,1,n-1,n\}}u_{1}(x)(1-f(x))^{2}+\mu\sum_{x\notin\{0,1,n-1,n\}}u_{1}(x)\left(\alpha(x)(1-f(1))^{2}+(1-\alpha(x))(1-f(n-1))^{2}\right)$
$+(1-\mu)\sum_{x\in\{0,1,n-1,n\}}u_{0}(x)f(x)^{2}+(1-\mu)\sum_{x\notin\{0,1,n-1,n\}}u_{0}(x)\left(\alpha(x)f(1)^{2}+(1-\alpha(x))f(n-1)^{2}\right)$ (Convexity)
$=\mu\sum_{x\in\{0,1,n-1,n\}}v^{\prime}_{1}(x)(1-f(x))^{2}+(1-\mu)\sum_{x\in\{0,1,n-1,n\}}v^{\prime}_{0}(x)f(x)^{2}$
where $u_{1}^{\prime}(0)=u_{1}(0),u_{1}^{\prime}(n)=u_{1}(n)$ and
$u_{1}^{\prime}(1)=u_{1}(1)+\sum_{2\leq x\leq n-2}\alpha(x)u_{1}(x)$
$u_{1}^{\prime}(n-1)=u_{1}(1)+\sum_{2\leq x\leq n-2}(1-\alpha(x))u_{1}(x)$
Thus we have
$\sum_{x}u_{1}^{\prime}(x)$ $=u_{1}^{\prime}(0)+u_{1}^{\prime}(1)+u_{1}^{\prime}(n-1)+u_{1}^{\prime}(n)$
$=u_{1}(0)+u_{1}(1)+u_{1}(n-1)+u_{1}(n)+\sum_{2\leq x\leq n-2}\alpha(x)u_{1}(x)+(1-\alpha)u_{1}(x)$
$=u_{1}(0)+u_{1}(1)+u_{1}(n-1)+u_{1}(n)+\sum_{2\leq x\leq n-2}u_{1}(x)$
$=1$
and
$\sum_{x}xu_{1}^{\prime}(x)$ $=u_{1}^{\prime}(1)+(n-1)u_{1}^{\prime}(n-1)+nu_{1}^{\prime}(n)$
$=u_{1}(1)+(n-1)u_{1}(n-1)+nu_{1}(n)+\sum_{2\leq x\leq n-2}\alpha(x)u_{1}(x)+(1-\alpha)(n-1)u_{1}(x)$
$=u_{1}(0)+u_{1}(1)+(n-1)u_{1}(n-1)+nu_{1}(n)+\sum_{2\leq x\leq n-2}xu_{1}(x)$
$=a$
Similarly we have $\sum_{x}u_{0}^{\prime}(x)=1,\sum_{x}xu_{0}^{\prime}(x)=b$. We can create a new information structure $\theta^{\prime}$ by setting $u_{1}^{\prime}(x)=\Pr_{\theta_{i}}[X=x|\omega=1]$ and $u_{0}^{\prime}(x)=\Pr_{\theta_{i}}[X=x|\omega=0]$. Notice that $\theta^{\prime}\in\Theta_{e}$. Thus for any $\theta\in\Theta$, we have $R(f,\theta)\leq R(f,\theta^{\prime})$. So
$R(\Theta)=\min_{f^{\prime}}R(f^{\prime},\Theta)\leq R(f,\Theta)\leq R(f,\Theta_{e})=R(f^{*},\Theta_{e})=R(\Theta_{e})$
However, since $\Theta_{e}\subset\Theta$, $R(\Theta_{e})\leq R(\Theta)$. So we obtain that $R(\Theta)=R(\Theta_{e})$.
Combining these two lemmas we only need to consider information structures with at most 4 possible reports and the reports are in $\{0,1,n-1,n\}$. In addition, $supp(u_{1})\leq 2,supp(u_{0})\leq 2$. There are at most 16 possible information structures and there exists an FPTAS for solving the finite number of information structures by guo2024algorithmic.
{lemma}
[guo2024algorithmic] Suppose $|\Theta|=n$, There exists an FPTAS which can find the $\epsilon$-optimal aggregator in $O(n/\epsilon)$.
If we apply the FPTAS in $\Theta_{e}$ we will find the $\epsilon$-optimal aggregator in $O(1/\epsilon)$, which completes our proof. The proofs are deferred to Section 4.1.
Regret
As we cannot obtain the closed-form of the optimal aggregator, we do not know the regret either. However, it is possible to estimate the regret, which is stated in section 4.2.2. The proof is deferred to Section 4.1.
{lemmarep}
Suppose $\Theta_{n}$ is the information structure with $n$ truthful experts. Then $R(\Theta_{n})=c-O(1/n)$ where $c$ is a constant related to $\mu,a,b$.
Proof.
Suppose $R_{n}$ is the regret when the number of experts is $n$. We want to estimate the difference $R_{n+1}-R_{n}$. Let $f_{1}=\arg\min_{f}\max_{\theta\in\Theta_{n}}R(f,\theta)$. As we prove in Section 4.2.2, we only need to consider point $0,1,n-1,n$. Then for any information structure $\theta_{1}$ in $\Theta_{n}$, we map it to another information structure $\theta_{2}\in\Theta_{n+1}$ by the following rules. If $\theta_{1}$ contains report $0,1$, let $\theta_{2}$ contains report $0,1$; if $\theta_{1}$ contains report $n-1$ or $n$, let $\theta_{2}$ contains report $n$ or $n+1$. Then the joint distribution is determined by the report space. Now we construct an aggregator $f_{2}$:
$f_{2}(x)=\left\{\begin{aligned} &f_{1}(x)&\quad x=0,1\\ &f_{1}(x-1)&\quad x=n,n+1\\ \end{aligned}\right.$ (3)
Then we have
$R(f_{2},\theta_{2})-R(f_{1},\theta_{1})$ $=\mathbb{E}_{x\sim\theta_{2}}[(f(x)-opt_{\theta_{2}}(x))^{2}]-\mathbb{E}_{x\sim\theta_{1}}[(f(x)-opt_{\theta_{1}}(x))^{2}]$
$=\sum_{x}\Pr_{\theta_{2}}[x](f(x)-opt_{\theta_{2}}(x))^{2}-\Pr_{\theta_{1}}[x](f(x)-opt_{\theta_{1}}(x))^{2}$
$\leq\sum_{x}\left|\Pr_{\theta_{2}}[x]-\Pr_{\theta_{1}}[x]\right|$ ($(f(x)-opt_{\theta_{2}}(x))^{2}\leq 1$)
$\leq O\left(\frac{1}{n(n+1)}\right)$
Thus $R(\Theta_{n+1})\leq R(f_{2},\Theta_{n+1})\leq R(f_{1},\Theta_{n})+O\left(\frac{1}{n(n+1)}\right)=R(\Theta_{n})+O\left(\frac{1}{n(n+1)}\right)$.
Add up all $n$ we have $R(\Theta_{n})\leq\sum_{k}O\left(\frac{1}{k(k+1)}\right)=c-O(\frac{1}{n})$
4.2.3 Adversarial Setting
Now we consider the adversarial setting. Surprisingly, when $\gamma$ is relatively small, the optimal aggregator has a closed-form, which is also a hard sigmoid function, with separation points in $k,n-k$. The difference between the L1 and L2 loss lies in the value of $f(k),f(n-k)$. We state the optimal aggregator in theorem 4.4.
Theorem 4.4.
When $0<\gamma<\min(\frac{a}{1+a},\frac{1-b}{2-b})$, the optimal aggregator is
$f^{*}(x)=\left\{\begin{aligned} &\frac{\mu(1-\gamma)(1-a)}{\mu(1-\gamma)(1-a)+(1-\mu)(1-2\gamma-(1-\gamma)b)}&\quad x\leq k\\ &\frac{\mu((1-\gamma)a-\gamma)}{\mu((1-\gamma)a-\gamma)+(1-\mu)(1-\gamma)b}&\quad x\geq n-k\\ &\frac{x-k}{n-2k}(f(n-k)-f(k))+f(k)&\quad otherwise\\ \end{aligned}\right.$ (4)
{toappendix}
proof of theorem 4.4.
When $k<(n-k)a,(n-k)b<n-2k$, we construct a bad information structure $\theta^{*},\sigma^{*}$ that $supp(u_{1})=\{k,n-k\},supp(u_{0})=\{0,n-2k\}$ and
$\sigma^{*}(x)=\left\{\begin{aligned} &k&\quad x=0,n-k\\ &0&\quad otherwise\\ \end{aligned}\right.$ (5)
Then we find the posterior
$\Pr_{\theta^{*},\sigma^{*}}[\omega=1|X=x]$ $=\frac{\Pr_{\theta^{*},\sigma^{*}}[\omega=1]\Pr_{\theta^{*},\sigma^{*}}[X=x|\omega=1]}{\Pr_{\theta^{*},\sigma^{*}}[\omega=1]\Pr_{\theta^{*},\sigma^{*}}[X=x|\omega=1]+\Pr_{\theta^{*},\sigma^{*}}[\omega=0]\Pr_{\theta^{*},\sigma^{*}}[X=x|\omega=0]}$
where
$\Pr_{\theta^{*},\sigma^{*}}[X=x|\omega=1]=\left\{\begin{aligned} &\frac{(n-k)a-k}{n-2k}&\quad x=n-k\\ &\frac{(n-k)(1-a)}{n-2k}&\quad x=k\\ &0&\quad otherwise\\ \end{aligned}\right.$ (6)
and
$\Pr_{\theta^{*},\sigma^{*}}[X=x|\omega=0]=\left\{\begin{aligned} &\frac{(n-k)b}{n-2k}&\quad x=n-2k\\ &\frac{n-2k-(n-k)b}{n-2k}&\quad x=0\\ &0&\quad otherwise\\ \end{aligned}\right.$ (7)
By simple calculate we find that $f^{*}(x)$ is the Bayesian posterior of $\theta^{*},\sigma^{*}$, thus $R(f,\theta^{*},\sigma^{*})\geq R(f^{*},\theta^{*},\sigma^{*})$ for any $f$. Now we want to prove that $R(f^{*},\theta^{*},\sigma^{*})\geq R(f^{*},\theta,\sigma)$ for any $\theta,\sigma$.
$\mathbb{E}_{\theta,\sigma}[\ell(f^{*})]$ $=\mu\sum_{x}\mathbb{E}_{\sigma}[v_{1}(x)(1-f^{*}(x+\sigma(x)))^{2}]+(1-\mu)\sum_{x}\mathbb{E}_{\sigma}[v_{0}(x)f^{*}(x+\sigma(x))^{2}]$
$\leq\mu\sum_{x}u_{1}(x)(1-f^{*}(x))^{2}+(1-\mu)\sum_{x}u_{0}(x)f^{*}(x+k)^{2}$ ($f^{*}$ is monotone)
$=\mu\sum_{x\geq k}u_{1}(x)(\alpha(x)k+(1-\alpha(x))n)(1-f^{*}(x))^{2}+\mu\sum_{x<k}u_{1}(x)(1-f^{*}(x))^{2}$ ($\alpha(x)=\frac{n-x}{n-k}$)
$+(1-\mu)\sum_{x\leq n-k}u_{0}(x)(\beta(x)0+(1-\beta(x))(n-k))f^{*}(x)^{2}+(1-\mu)\sum_{x>n-k}u^{\prime}_{0}(x)f^{*}(x)^{2}$ ($\beta(x)=\frac{n-k-x}{n-k}$)
$\leq\mu\sum_{x\in\{k,n\}}u^{\prime}_{1}(x)(1-f^{*}(x))^{2}+\mu\sum_{x<k}u_{1}^{\prime}(x)(1-f^{*}(x))^{2}$ ($(1-f^{*})^{2}$ is convex in $[k,n]$)
$+(1-\mu)\sum_{x\in\{0,n-k\}}u^{\prime}_{0}(x)f^{*}(x)^{2}+(1-\mu)\sum_{x>n-k}u^{\prime}_{0}(x)f^{*}(x)^{2}$ (${f^{*}}^{2}$ is convex in $[0,n-k]$)
Notice that $u_{1}^{\prime}(x)$ and $u_{0}^{\prime}(x)$ are still feasible distribution vectors according to calculations in previous proof.
Now we calculate the maximum of
$\sum_{x\in\{k,n\}}u_{1}^{\prime}(x)(1-f^{*}(x))^{2}+\sum_{x<k}u_{1}^{\prime}(x)(1-f^{*}(x))^{2}$
subject to
$\sum_{x}u_{1}^{\prime}(x)=1$
$\sum_{x}xu_{1}^{\prime}(x)=(n-k)a$
Notice that $f^{*}(x)$ is constant when $x\leq k$, for any $x<k$, we write $\alpha(x)=\frac{n-x}{n-k}$, then
$u_{1}^{\prime}(x)(1-f^{*}(x))^{2}$ $=u_{1}^{\prime}(x)(\alpha(x)(1-f^{*}(x))^{2}+(1-\alpha)(1-f^{*}(x))^{2}$
$\leq u_{1}^{\prime}(x)(\alpha(x)(1-f^{*}(k))^{2}+(1-\alpha)(1-f^{*}(n))^{2}$ ($f^{*}(x)=f^{*}(k)\leq f^{*}(n)$)
Thus we can replace any $x<k$ with the combination of $k,n$ until there are only reports $k,n$ left. That is,
$\sum_{x\in\{k,n\}}u_{1}^{\prime}(x)(1-f^{*}(x))^{2}+\sum_{x<k}u_{1}^{\prime}(x)(1-f^{*}(x))^{2}$
$\leq$ $\sum_{x=\{k,n\}}u^{\prime\prime}_{1}(x)(1-f^{*}(x))^{2}$
where
$\sum_{x}u_{1}^{\prime\prime}(x)=1$
$\sum_{x}xu_{1}^{\prime\prime}(x)=(n-k)a$
Similarly, we have
$\sum_{x\in\{0,n-k\}}u^{\prime}_{0}(x)f^{*}(x)^{2}+(1-\mu)\sum_{x>n-k}u^{\prime}_{0}(x)f^{*}(x)^{2}$
$\leq$ $\sum_{x=\{0,n-k\}}u^{\prime\prime}_{0}(x)f^{*}(x)^{2}$
where
$\sum_{x}u_{0}^{\prime\prime}(x)=1$
$\sum_{x}xu_{0}^{\prime\prime}(x)=(n-k)b$
In fact, $u_{1}^{\prime\prime},u_{0}^{\prime\prime}$ is the same as $\theta$. So $R(f^{*},\theta,\sigma)\leq R(f^{*},\Theta^{*},\sigma^{*})$, which complete our proof. ∎
We can also obtain the value of regret.
Corollary 4.5.
When $0<\gamma<\min(\frac{a}{1+a},\frac{1-b}{2-b})$, the regret $R(\Theta,\Sigma)$ is
$\frac{(\mu-1)\mu(\gamma-1)\left(\gamma\left(-a^{2}\mu+(b-2)b(\mu-1)+\mu\right)+(\mu(a-b)(a+b-1)+(b-1)b)\right)}{(-(a+1)\gamma\mu+a\mu+b(\mu-1)(\gamma-1))(\gamma(-(a+1)\mu+b(\mu-1)+2)+(a\mu-b\mu+b-1))}.$
Proof Sketch
First, we create a bad information structure with $supp(v_{1}),supp(v_{0})\subset\{k,n-k\}$. Then for any aggregator, the regret will be larger than the regret in corollary 4.5.
Next, we provide a hard sigmoid function as stated in theorem 4.4. We prove that for any information structures, the regret under this hard sigmoid function will be less than the regret in corollary 4.5. The key idea here is to decompose the information structures into several “simple” information structures and use the convexity of regret function. The proof is deferred to Section 4.1.
Discussion
The adversarial setting has some essential differences from the non-adversarial setting. In the non-adversarial setting, the regret is related to the number of experts $n$. Thus it is not simply the special case $\gamma=0$ in the adversarial setting, whose regret is a constant when $\gamma=0$. The reason is that there is a trade-off between the performance of the optimal aggregator and the benchmark. If we follow the same flow of adversarial setting and create a bad information structure such that $supp(v_{1}),supp(v_{0})\subset\{0,n\}$ in the non-adversarial setting, the loss of benchmark is not zero, although the loss of any aggregator will be bad in this information structure. On the other hand, as for the information structures with $supp(v_{1}),supp(v_{0})\subset\{0,1,n-1,n\}$, the loss of benchmark could be zero, while the loss of aggregator will also decrease. Notice that the final regret is $loss(f)-loss(opt)$. Thus, we cannot determine which information structure is the worst for any aggregator in the non-adversarial setting easily.
4.3 Comparison for Different Aggregators
Figure 2 summarizes the optimal aggregators under different settings. The aggregators under L2 loss are more conservative compared to L1 loss because it is closer to the middle point 0.5. Compared to the non-adversarial setting, the adversarial aggregators have a higher gradient in the middle case ($k\leq x\leq n-k$), while they treat the corner case ($x\geq n-k$ and $x\leq k$) indifferently.
Figure 3 shows the regret $R(\Theta,\Sigma)$ for different number of adversaries under L1 loss. For a small group of adversaries, the regret is quite stable and increases asymptotically linearly. However, when the ratio of adversaries $\gamma$ grows over $1/4$, the regret increases rapidly. When $\gamma$ approximates $1/2$, the optimal aggregators do not satisfy constraints in theorem 4.2,theorem 4.4. Thus the regret curve will be different from the figure. The regret will reaches the maximum $1/2$ (L1 loss) and $1/4$ (L2 loss) when $\gamma\approx 1/2$.
0011112222333344445555666677778888999910101010000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111Number of Experts Reporting HOutput of AggregatorsOptimal AggregatorNon-adversarial: Optimal Aggregator, L1 lossAdversarial: Optimal Aggregator, L1 lossAdversarial: Optimal Aggregator, L2 lossNon-adversarial: Optimal Aggregator, L2 loss
Figure 2: Comparison for optimal aggregators under different settings. The x-axis is the number of experts who report ”H” and the y-axis is the output of optimal aggregators. We set $n=10,k=2,a=0.8,b=0.1,\mu=0.5$. The dashed lines are the non-adversarial aggregators, and the solid lines are the adversarial aggregators. The red lines are aggregators under L2 loss while the blue lines are under L1 loss.
(a) L1 loss
(b) L2 loss
Figure 3: Regret for the different number of adversaries under different loss functions. The x-axis is the number of adversaries and the y-axis is the value of regret. We set $n=100,a=0.8,b=0.1,\mu=0.5$.
5 Numerical Experiment
In this section, we will evaluate our aggregators in a machine learning application, which is also known as ensemble learning. In ensemble learning, we have several basic but different models. Every model will give its prediction about the tasks. We want to aggregate their predictions to obtain a better model with higher accuracy. There also exist data poisoning attacks (jagielski2018manipulating) in the ensemble learning process in practice, which corresponds to the adversarial setting.
5.1 Setup
Now we introduce how to apply our decision aggregation framework in the ensemble learning setting. We consider a classification problem.
  • World State $\omega$: it is the true class $y_{j}$ of the data point $d_{j}$.
  • Expert $i$: it is a black-box machine learning model $M_{i}$, which will take the data point $d_{j}$ as the input and output a prediction for its class $M_{i}(d_{j})$.
  • Signal $s_{i}$: suppose we have a training dataset $\mathcal{D}$. For each model $M_{i}$, the signal $s_{i}$ is defined by its training dataset $\mathcal{D}_{i}\subset\mathcal{D}$.
  • Report $x_{i}$: it is model $M_{i}$’s output class. We do not consider the confidence of models.
  • Benchmark Function $opt$: it is defined by the best model trained using the full dataset $\mathcal{D}$.
In our experiment, we select the cifar-10 dataset, which contains images of 10 different classes. To apply our binary signals setting, we need some adjustments to convert the task into a binary decision problem. That is, whether the image belongs to a special class (e.g. cat) or not. To ensure the symmetric assumption and keep the diversity of the models, we train 100 models using the same machine learning backbone according to David2018. For each model $M_{i}$, we construct the dataset $\mathcal{D}_{i}$ by uniformly sampling 20000 images from the original training dataset, which contains 50000 images. We train 10 epochs by GPU RTX 3060 and CPU Intel i5-12400. The average accuracy of models is around 85%.
Estimation For Parameters
To apply our aggregator in theorem 4.4, we need three important parameters, the prior $\mu$, the probability $a$ of yes answer when the ground truth is yes, and the probability $b$ of yes answer when the ground truth is no. In practice, it is impossible to know the true value due to the imperfect knowledge of the data distribution. Fortunately, we can estimate them by the empirical performance of models in the training set. If the training set is unbiased samples of the true distribution, the empirical estimator is also unbiased for the true parameters.
Adversary Models
We test two different kinds of adversarial strategies. First is the extreme strategy. That is, the adversaries always report the opposite forecast to the majority of truthful experts. Second is the random strategy, which will randomly report yes or no with equal probability.
Aggregators
We test three aggregators: the majority vote, which picks the answers of the majority of experts; the averaging, which outputs the ratio of models answering “yes”; and the optimal aggregator calculated as in previous sections.
5.2 Results
We test the performance of different aggregators under a diverse number of adversaries and Figure 4 shows our results for L2 loss. L1 loss case is deferred to Section 5.2. Our optimal aggregator outperforms other aggregators in any situation. Even though the majority has a close performance to the optimal aggregator, which means it is a good approximation in the ensemble learning setting. This is mainly due to the high accuracy of each model. When the group of adversaries is small, the averaging performs well. However, when the group becomes large, the averaging quickly collapses into a poor choice. Thus averaging is very unreliable in the adversarial setting.
When the adversaries use the random strategy, for all the aggregators, their performance is better than the extreme strategy. In particular, the majority vote will not be affected by the adversaries in expectation since it will not change the majority. When the adversaries are small, the averaging even outperforms the optimal aggregator. It may result in the conservative tendency of the optimal aggregator.
(a) Extreme Strategy
(b) Random Strategy
Figure 4: The Performance of Different Aggregators Under Different Adversarial Strategies: The x-axis is the number of adversaries we add. The number of truthful experts is 100. The y-axis is the regret.
{toappendix}
[Experiment]
(a) Extreme Strategy
(b) Random Strategy
Figure 5: The Performance of Different Aggregators Under Different Adversarial Strategies: The x-axis is the number of adversaries we added to experts. The number of truthful experts is 100. The y-axis is the regret.
Figure 5 shows the results under L1 loss. The optimal aggregator and majority vote have close performance. It also reflects our discussion that L1 loss will encourage the aggregators to output a decision, which is beneficial for the majority vote.
6 Extension
In this section, we extend our adversarial information aggregation problem into a general case. In the general setting, we can capture a non-binary state space $\Omega$, other families of information structures $\Theta$, and other kinds of reports $x_{i}$ such as the posterior forecast $x_{i}=\Pr_{\theta}[\omega|s_{i}]$.
6.1 Problem Statement
Suppose the world has a state $\omega\in\Omega$, $|\Omega|=m$. There are $n$ experts and each expert $i$ receives a private signal $s_{i}$ in a signal space $\mathcal{S}_{i}$. Let $\mathcal{S}=\mathcal{S}_{1}\times\mathcal{S}_{2}\times...\times\mathcal{S}_{n}$. They are asked to give a report from a feasible choice set $X$ according to their private signals. We denote the joint distribution, or information structure, over the state and signals by $\theta\in\Delta_{\Omega\times\mathcal{S}}$. For simplicity, we denote ${\bm{x}}_{T}=(x_{t})_{t\in T}$ as the sub-vector for any vector ${\bm{x}}$ and index set $T$.
We assume the experts are either truthful or adversarial. Let $T$ denote the set of truthful experts who will give their best report truthfully. $A$ is the set of adversarial experts who will collude and follow a randomized strategy $\sigma_{A}:\mathcal{S}_{A}\to\Delta_{X^{k}}$ depending on their private signals, where $k=\gamma n$ is the number of adversarial experts. The family of all possible strategies is $\Sigma$. We assume $\gamma<1/2$ such that there exist at least half truthful experts, otherwise no aggregators can be effective.
Notice that the ability of adversarial experts can be modeled by their private signals. For example, the omniscient adversarial experts can know truthful experts’ signals $\mathcal{S}_{R}$ and the world’s true state $\omega$. The ignorant adversarial experts are non-informative, i.e. receiving nothing.
The aggregator is an anonymous function $f(\cdot)\in\mathcal{F}$ which maps ${\bm{x}}\in X^{n}$ to a distribution ${\bm{y}}\in\Delta_{\Omega}$ over the world state. We define a loss function $\ell({\bm{y}},\omega):\Delta_{\Omega}\times\Omega\to R^{+}$, indicating the loss suffered by the aggregator when the aggregator’s predicted distribution of the state is ${\bm{y}}$ and the true state is $\omega$. The expected loss is $\mathbb{E}_{\theta,\sigma}[\ell(f({\bm{x}}),\omega)]$. We assume $\ell$ is symmetric and convex for any state, which means we can abbreviate $\ell({\bm{y}},\omega)$ by $\ell(y_{\omega})$. Without loss of generality, we assume $\ell(0)=1,\ell(1)=0$ and $\ell(\cdot)$ is decreasing. In particular, we consider the L1 loss (or absolute loss) $\ell_{1}(y)=1-y$ and the L2 loss (or square loss), $\ell_{2}(y)=(1-y)^{2}$. We define a benchmark function, that gives the optimal result given the joint distribution and truthful experts’ reports:
$opt_{\theta}({\bm{x}}_{T})=\operatorname*{arg\,min}_{f^{\prime}\in\mathcal{F}}\mathbb{E}_{\theta}[\ell(f^{\prime}({\bm{x}}_{T})_{\omega})]$
to minimize the expected loss.
Under the L1 loss, $opt_{\theta}({\bm{x}}_{T})_{\omega}=\mathbbm{1}(\omega=\arg\max_{\omega}\Pr_{\theta}[\omega|{\bm{x}}_{T}])$, which is the maximum likelihood. Under the L2 loss, $opt_{\theta}({\bm{x}}_{T})_{\omega}=\Pr_{\theta}[\omega|{\bm{x}}_{T}]$, which is the Bayesian posterior.
Regret Robust Paradigm
Given a family of joint distributions $\Theta$ and a family of strategies $\Sigma$, a set of aggregators $\mathcal{F}$, we aim to minimize the expected loss in the worst information structure. That is, we want to find an optimal function $f^{*}$ to solve the following min-max problem:
$R(\Theta,\Sigma)=\inf_{f\in\mathcal{F}}\sup_{\theta\in\theta,\sigma\in\Sigma}\mathbb{E}_{\theta,\sigma}[\ell(f({\bm{x}})_{\omega})]-\mathbb{E}_{\theta}[\ell(opt_{\theta}({\bm{x}}_{T})_{\omega})].$
Again we define $R(f,\theta,\sigma)=\mathbb{E}_{\theta,\sigma}[\ell(f({\bm{x}})_{\omega})]-\mathbb{E}_{\theta}[\ell(opt_{\theta}({\bm{x}}_{T})_{\omega})]$ and $R(f,\Theta,\Sigma)=\sup_{\theta\in\Theta,\sigma\in\Sigma}R(f,\theta,\sigma)$ for short.
6.2 Negative Results For General Model
{toappendix}
We first provide an auxiliary lemma to characterize the behavior of adversaries. {lemma} Assume $\Sigma_{p}=\{\sigma|\sigma\in\Sigma\ and\ \sigma\ is\ a\ pure\ strategy\}$. Then $R(f,\Theta,\Sigma)=R(f,\Theta,\Sigma_{p})$ for any $f$ and $\Theta$.
Proof.
On the one hand, $\Sigma_{p}\in\Sigma$, so $R(f,\theta,\Sigma)\geq R(f,\theta,\Sigma_{p})$.
On the other hand, for any $f,\theta,\sigma$,
$R(f,\theta,\sigma)$ $=\mathbb{E}_{\theta}\mathbb{E}_{\sigma}[\ell(f({\bm{x}}_{T},\sigma({\bm{s}}_{A})),\omega)]-\mathbb{E}_{\theta}[\ell(opt_{\theta}({\bm{x}}_{T})_{\omega})]$
$\leq\mathbb{E}_{\theta}[\max_{{\bm{x}}_{A}^{\prime}}\ell(f({\bm{x}}_{T},{\bm{x}}_{A}^{\prime}({\bm{s}}_{A})),\omega)]-\mathbb{E}_{\theta}[\ell(opt_{\theta}({\bm{x}}_{T})_{\omega})]$
Fix the aggregator $f$ and distribution $\theta$, we can let $\sigma^{\prime}({\bm{s}}_{A})=\arg\max_{{\bm{x}}_{A}^{\prime}}\mathbb{E}_{\theta}[\ell(f({\bm{x}}_{T},{\bm{x}}_{A}^{\prime}({\bm{s}}_{A})),\omega)]$ (We arbitrarily select one ${\bm{x}}_{A}^{\prime}$ when there is multiple choices). Since $\sigma^{\prime}$ is a pure strategy, $\sigma^{\prime}\in\Sigma^{d}$. For any $\sigma\in\Sigma$, $R(f,\theta,\sigma^{\prime})\geq R(f,\theta,\sigma)$. Thus $R(f,\theta,\Sigma)\leq R(f,\theta,\Sigma_{p})$, which completes our proof. ∎
6.3 Extension to Multi-State Case
For the multi-state case: $|\Omega|>2$, we can also define the sensitive parameter by enumerating each state:
Definition 6.1 (sensitive parameter, multiple).
When $|\Omega|>2$, for any benchmark function $opt$, the sensitive parameter is defined by
$S(opt,k)=\max_{\theta,\theta^{\prime}\in\Theta,d({\bm{x}}_{T},{\bm{x}}_{T}^{\prime})\leq k,\omega\in\Omega}|opt_{\theta}({\bm{x}}_{T})_{\omega}-opt_{\theta^{\prime}}({\bm{x}}_{T}^{\prime})_{\omega}|.$
First, we prove that for the general model, aggregators are vulnerable to adversaries. A direct observation is that fixing the number of experts $n$ while increasing the number of adversaries $k$ will not decrease the regret $R(\Theta,\Sigma)$. Since the new adversaries can always pretend to be truthful experts. In most cases, when $\gamma\approx 1/2$, it is impossible to design any informative aggregator since an adversarial expert can report opposite views to another truthful expert.
One natural question is how many adversaries we need to attack the aggregator. For this question, we provide a negative result. The following lemma shows that for a large family of information structures, a few adversaries are enough to fool the aggregator. Surprisingly, the number of adversaries is independent of the number of truthful experts but grows linearly with the number of states. In special, in the binary state setting, one adversary is enough to completely fool the aggregator. The proof is deferred to Section 6.2.
{lemmarep}
Assume the truthful experts are asked to report the posterior $x_{i}(\omega)=\Pr_{\theta}[\omega|s_{i}]$. We define the fully informative expert who always knows the true state and the non-informative expert who only knows the prior. If $\Theta$ includes all information structures consisting of these two types of experts, then for $k\geq m-1$, the optimal aggregator is the uniform prediction. That is, $f^{*}=(\frac{1}{m},\cdots,\frac{1}{m})$ and $R(\theta,\sigma)=\ell(\frac{1}{m})$.
{appendixproof}
We select one fully informative expert, which means she will report a unit vector $\bm{e}_{i}=(0,\cdots,1,0,\cdots,0)$ with $e_{i}=1$ and $e_{j}=0$ for any $j\neq i$. Other experts are non-informative and will always report the uniform prior $(\frac{1}{m},\cdots,\frac{1}{m})$. Then let the $m-1$ adversaries report other unit vectors $\bm{e}_{j}$ for $i\neq j$. Other adversaries also report the uniform prior. In that case, the aggregator will always see the same reports ${\bm{x}}^{0}$, and the benchmark can follow the informative expert and suffer zero loss. Since $\ell$ is convex and $\sum_{\omega\in\Omega}f({\bm{x}}^{0})_{\omega}=1$, then for any $f$, $R(f,\theta,\sigma)=\sum_{\omega\in\Omega}\ell(f({\bm{x}}^{0})_{\omega})\geq\ell(\frac{1}{m})$. Thus $R(\Theta,\Sigma)\geq\ell(\frac{1}{m})$.
On the other hand, when $f^{*}=(\frac{1}{m},\cdots,\frac{1}{m})$, $R(f^{*},\theta,g)=\sum_{\omega\in\Omega}\Pr[\omega]\ell(\frac{1}{m})=\ell(\frac{1}{m})$ for any $\theta,\sigma$. So the uniform prediction is the optimal aggregator and $R(\Theta,\Sigma)=\ell(\frac{1}{m})$.
The uniform prediction means that we cannot obtain any additional information from reports. It is almost impossible in the non-adversarial setting. We provide a common setting as an example.
Example 6.2 (Conditionally Independent Setting).
[…]
Corollary 6.3.
In the conditionally independent setting, if we select $\ell$ as the L2 loss, then for any $n$ and $k\geq 1$, $R=1/4$.
Notice that in the non-adversarial setting, when $n\rightarrow\infty,R\rightarrow 1/4$ (arieli2018robust), which is a strict condition. However, in the adversarial setting, we only use one adversary to obtain the same bad regret.
6.4 Estimate the Regret $R(\Theta,\Sigma)$
We have provided a negative result when there are enough adversarial experts. In this section, we want to extend our result regarding the regret to any number of adversarial experts setting. It will give us a further understanding of the effect of adversaries.
Intuitively, when there exist important experts, it is easier to disturb the aggregator because adversaries can always imitate an important expert but hold the opposite view. In other words, the regret will increase because the DM is non-informative while the benchmark can predict accurately. Thus we can use the importance of an expert to bound the regret. The remaining question is, how to quantify the importance of an expert? We find that the benchmark function is a proper choice. We state our main result as theorem 6.4.
Theorem 6.4.
If $\Theta$ is $\alpha$-regular as defined in definition 6.8, for L2 loss function $\ell_{2}$, there exists a function $S(opt,k)$ depending on the benchmark function $opt$ and number of adversaries $k$ such that
$S(opt,k)^{2}\geq R(\Theta,\Sigma)\geq\frac{\alpha}{4}S(opt,k)^{2}.$
The theorem shows that we can only use the benchmark function to design a metric for the difficulty of adversarial information aggregation. Now we give the formula of $S(\cdot)$. First, we define the distance of reports.
Definition 6.5 (distance of reports).
For any vector ${\bm{x}}$, we can define its histogram function $h_{{\bm{x}}}(x)=\sum_{i}\mathbbm{1}(x_{i}=x)$. For any pair of reports ${\bm{x}}_{1},{\bm{x}}_{2}$ with the same size, their distance is defined by the total variation distance between $h_{{\bm{x}}_{1}}(x)$ and $h_{{\bm{x}}_{2}}(x)$.
$d({\bm{x}}_{1},{\bm{x}}_{2})=TVD(h_{{\bm{x}}_{1}},h_{{\bm{x}}_{2}})=1/2\sum_{x}\left|h_{{\bm{x}}_{1}}(x)-h_{{\bm{x}}_{1}}(x)\right|.$
If $d({\bm{x}}_{1},{\bm{x}}_{2})=0$, we say that ${\bm{x}}_{1}$ and ${\bm{x}}_{2}$ are indistinguishable, which means they only differ in the order of reports. In fact, $d({\bm{x}}_{1},{\bm{x}}_{2})$ is the minimal number of adversaries needed to make two reports indistinguishable. We also denote $d({\bm{x}})=\sum_{x}h_{\bm{x}}(x)$.
{lemmarep}
Suppose the truthful experts will report ${\bm{x}}_{1}\in X^{n-k}$ or ${\bm{x}}_{2}\in X^{n-k}$. If and only if $k\geq d({\bm{x}}_{1},{\bm{x}}_{2})$, there exists ${\bm{x}}_{A}^{1},{\bm{x}}_{A}^{2}\in X^{k}$ such that $d(({\bm{x}}_{1},{\bm{x}}_{A}^{1}),({\bm{x}}_{2},{\bm{x}}_{A}^{2}))=0$. That is, $d({\bm{x}}_{1},{\bm{x}}_{2})$ is the minimal number of adversaries needed to ensure the aggregator sees the same report vector ${\bm{x}}$ in these two cases.
{appendixproof}
Consider the final report vector ${\bm{x}}$ seen by the aggregator, we have $h_{{\bm{x}}}(x)\geq\max(h_{{\bm{x}}_{1}}(x),h_{{\bm{x}}_{2}}(x))$ for any $x$. Thus to convert ${\bm{x}}_{2}$ to ${\bm{x}}$, we need at least
$\sum_{x}\lvert h_{{\bm{x}}}(x)-h_{{\bm{x}}_{2}}(x)\rvert$ $\geq\sum_{x}\max(h_{{\bm{x}}_{1}}(x)-h_{{\bm{x}}_{2}}(x),0)$
$=1/2\sum_{x}\left(\lvert h_{{\bm{x}}_{1}}(x)-h_{{\bm{x}}_{2}}(x)\rvert+h_{{\bm{x}}_{1}}(x)-h_{{\bm{x}}_{2}}(x)\right)$
$=1/2\sum_{x}\lvert h_{{\bm{x}}_{1}}(x)-h_{{\bm{x}}_{2}}(x)\rvert$ ($\sum_{x}h_{{\bm{x}}_{1}}(x)=\sum_{x}h_{{\bm{x}}_{2}}(x)=n$)
Similarly, to convert ${\bm{x}}_{1}$ to ${\bm{x}}$, we need $\lvert h_{{\bm{x}}_{1}}(x)-h_{{\bm{x}}_{2}}(x)\rvert$ adversaries, which complete our proof.
The proof is deferred to Section 6.2. Now we define the sensitive parameter in the binary state case, where the benchmark function can be represented by a real number. Intuitively, it measures the greatest change $k$ experts can make regarding the benchmark function.
Definition 6.6 (sensitive parameter, binary).
When $|\Omega|=2$, for any benchmark function $opt$, the sensitive parameter is defined by
$S(opt,k)=\max_{\theta,\theta^{\prime}\in\Theta,d({\bm{x}}_{T},{\bm{x}}_{T}^{\prime})\leq k}|opt_{\theta}({\bm{x}}_{T})-opt_{\theta^{\prime}}({\bm{x}}_{T}^{\prime})|$
We define the sensitive parameter in the binary state case. However, it is easy to generalize it to multi-state cases. We will discuss it in section 6.2.
Proof Sketch of Theorem 6.4
We provide a proof sketch for theorem 6.4 and the proofs are deferred to Section 6.2. First, we consider the upper bound. We construct an aggregator and prove its regret can be upper bounded by the sensitive parameter (Section 6.4). {lemmarep} For any benchmark function $opt$,
$R(\Theta,\Sigma)\leq\ell(S(opt,k))$
{appendixproof}
First, we construct a naive aggregator in the binary setting.
Definition 6.7 (naive aggregator).
We use the average of the maximum and minimal prediction over all possible situations as the naive aggregator. Formally, we define $u({\bm{x}})=\max_{\theta}opt_{\theta}({\bm{x}})$ and $l({\bm{x}})=\min_{\theta}opt_{\theta}({\bm{x}})$ for any ${\bm{x}}\in X^{n-k}$. The aggregator
$f^{na}({\bm{x}})=1/2\max_{d({\bm{x}}^{\prime})=n-k,d({\bm{x}},{\bm{x}}^{\prime})=k}u({\bm{x}}^{\prime})+1/2\min_{d({\bm{x}}^{\prime})=n-k,d({\bm{x}},{\bm{x}}^{\prime})=k}l({\bm{x}}^{\prime})$
By the definition of $S(opt,k)$, for any $\theta\in\Theta,\sigma\in\Sigma$,
$R(\theta,\sigma)$ $\leq R(f^{na},\theta,\sigma)$
$\leq\sup_{\theta\in\Theta,\sigma\in\Sigma}\mathbb{E}_{\theta,\sigma}[\ell(f^{na}({\bm{x}}),\omega)]-\mathbb{E}_{\theta}[\ell(opt_{\theta}({\bm{x}}_{T}),\omega)]$
$=\sup_{\theta\in\Theta,\sigma\in\Sigma}\mathbb{E}_{\theta,\sigma}[\ell(f^{na}({\bm{x}}),\omega)-\ell(opt_{\theta}({\bm{x}}_{T}),\omega)]$
$\leq\max_{\theta\in\Theta,\sigma\in\Sigma}\ell(|f^{na}({\bm{x}})-opt_{\theta}({\bm{x}}_{T})|)$
$\leq\ell(S(opt,k)/2)$
However, The proof fails in the multi-state case since $f^{na}({\bm{x}})$ is not a distribution, thus illegal in this setting. Instead, we need another aggregator.
Suppose $u({\bm{x}})_{\omega}=\max_{\theta}opt_{\theta}({\bm{x}})_{\omega}$. Then it is obvious that $\sum_{\omega}u({\bm{x}})_{\omega}\geq\sum_{\omega}opt_{\theta}({\bm{x}})_{\omega}=1$. Similarly we have $l({\bm{x}})_{\omega}=\min_{\theta}opt_{\theta}({\bm{x}})_{\omega}$ and $\sum_{\omega}l({\bm{x}})_{\omega}\leq\sum_{\omega}opt_{\theta}({\bm{x}})_{\omega}=1$.
Thus for every ${\bm{x}}$ there exists a vector $f^{\prime}({\bm{x}})$ such that for any $\omega\in\Omega$, $l({\bm{x}})_{\omega}\leq f^{\prime}({\bm{x}})_{\omega}\leq u({\bm{x}})_{\omega}$ and $\sum_{\omega}f({\bm{x}})_{\omega}=1$. We can pick $f^{\prime}({\bm{x}})$ as the aggregator and for any $\theta\in\Theta,\sigma\in\Sigma$,
$R(\theta,\sigma)$ $\leq R(f^{\prime},\theta,\sigma)$
$\leq\sup_{\theta\in\Theta,\sigma\in\Sigma}\mathbb{E}_{\theta,\sigma}[\ell(f^{\prime}({\bm{x}}),\omega)]-\mathbb{E}_{\theta}[\ell(opt_{\theta}({\bm{x}}_{T}),\omega)]$
$=\sup_{\theta\in\Theta,\sigma\in\Sigma}\mathbb{E}_{\theta,\sigma}[\ell(f^{\prime}({\bm{x}}),\omega)-\ell(opt_{\theta}({\bm{x}}_{T}),\omega)]$
$\leq\max_{\theta\in\Theta,\sigma\in\Sigma}\ell(|f^{\prime}({\bm{x}})-opt_{\theta}({\bm{x}}_{T})|)$
$\leq\ell(S(opt,k))$
Now we consider the lower bound of $R(\Theta,\Sigma)$ (Section 6.4). A basic idea is that we can construct a coupling of two information structures whose reports have a distance smaller than $k$. Then the adversaries can make the coupling indistinguishable. However, if the probability of the worst case is too small, then their contribution to the regret is also negligible. Thus we need to constrain the information structures (Definition 6.8).
Definition 6.8 ($\alpha$-regular information structure).
An information structure $\theta$ is $\alpha$-regular if every possible report vector ${\bm{x}}$ will appear with probability at least $\alpha$:
$\min_{{\bm{x}}:\Pr_{\theta}[{\bm{x}}]>0}\Pr_{\theta}[{\bm{x}}]>\alpha.$
{lemmarep}
If $\theta$ is $\alpha$-regular, for any benchmark function $opt$ and L2 loss $\ell_{2}$, $R(\Theta,\Sigma)\geq\frac{\alpha}{4}S(opt,k)^{2}$.
{appendixproof}
Fix $\alpha$, let the maximum of $S(opt,k)$ is obtained by $\theta,\theta^{\prime},{\bm{x}}_{T},{\bm{x}}_{T}^{\prime}$. We construct an information structure as follows. Suppose the joint distribution $\theta^{\prime\prime}$ is the mixture of $\theta$ and $\theta^{\prime}$ with equal probability. By the definition of $d({\bm{x}}_{T},{\bm{x}}_{T}^{\prime})$, there exists $\sigma^{\prime}$ such that $h_{{\bm{x}}_{T},\sigma^{\prime}({\bm{x}}_{T})}=h_{{\bm{x}}_{T}^{\prime},\sigma^{\prime}({\bm{x}}_{T}^{\prime})}$. Then we have
$R(\theta,\sigma)$ $\geq R(\theta^{\prime\prime},\sigma^{\prime})$
$\geq\min_{f}\mathbb{E}_{\theta^{\prime\prime},\sigma^{\prime}}[\ell(f({\bm{x}}),\omega)]-\mathbb{E}_{\theta^{\prime\prime}}[\ell(opt_{\theta}^{\prime\prime}({\bm{x}}_{T}^{\prime\prime}),\omega)]$
$=\min_{f}\mathbb{E}_{\theta^{\prime\prime},g^{\prime}}[(f({\bm{x}})-opt_{\theta^{\prime\prime}}({\bm{x}}_{T}^{\prime\prime}))^{2}]$
$=1/2\min_{f}\Pr_{\theta}[{\bm{x}}_{T}](f({\bm{x}})-opt_{\theta}({\bm{x}}_{T}))^{2}+\Pr_{\theta^{\prime}}[{\bm{x}}_{T}^{\prime}](f({\bm{x}})-opt_{\theta^{\prime}}({\bm{x}}_{T}^{\prime}))^{2}$
$\geq\frac{\Pr_{\theta}[{\bm{x}}_{T}]\Pr_{\theta^{\prime}}[{\bm{x}}_{T}^{\prime}]}{2(\Pr_{\theta}[{\bm{x}}_{T}]+\Pr_{\theta^{\prime}}[{\bm{x}}_{T}^{\prime}])}(opt_{\theta}({\bm{x}}_{T})-opt_{\theta^{\prime}}({\bm{x}}_{T}^{\prime}))^{2}$
$\geq\frac{\alpha}{4}S(opt,k)^{2}$
Combining section 6.4 and section 6.4 we infer theorem 6.4. We discuss the value of $S(opt,k)$ in some common settings in Section 6.2.
{toappendix}
6.5 Case Studies
To study the performance of our bound, we discuss the value of $S(opt,k)$ in several common settings. Our results show that for many information structures, the upper bound is tight.
Conditional Independent Setting
In the conditional independent setting we introduced before, it is obvious that $S(opt,k)=1$ for any $k\geq 1$. By corollary 6.3 the bound in section 6.4 is tight.
Linear Setting
We call a joint distribution is linear if $opt_{\theta}({\bm{x}})=\beta\mu+(1-\beta)\sum_{i}t_{i}/n$ for some constants $\beta\in(0,1)$, prior $\mu\in(0,1)$, and reports $x_{i}=\beta\mu+(1-\beta)t_{i}$ for some private signals $t_{i}\in[0,1]$. An example of the linear setting is that the experts will observe the same number of samples from a binary distribution and they share a common prior $\mu$. We want to estimate the positive probability of the random variable.
In this case, $S(opt,k)=O(\frac{k}{n})$. Thus $R(\Theta,\Sigma)=O(\gamma)$.
On the other hand, as for the lower bound of $S(opt,k)$, we can prove that for any information structures, $S(opt,k)$ increases with linear complexity of $\gamma$. Therefore, in the linear setting, $S(opt,k)=\Theta(\gamma)$.
{lemma}
For any $opt$, $S(opt,k)/S(opt,n)=\Omega(\gamma)$ when $n$ is large enough.
Proof.
First, it is obvious that $S(opt,k)$ is increasing.
Suppose $S(opt,n)$ is obtained by $\theta,{\bm{x}},\theta^{\prime},{\bm{x}}^{\prime}$. Then we can construct a sequence $\{{\bm{x}}_{t}\}_{t\in[T]}$ such that for every $t\in[T-1]$, $d({\bm{x}}_{t},{\bm{x}}_{t+1})=k$. Moreover, $d({\bm{x}},{\bm{x}}_{1})=k$,$d({\bm{x}}_{T},{\bm{x}}^{\prime})\leq k$. Since $d({\bm{x}},{\bm{x}}^{\prime})=n$, we have $T\leq\frac{n}{k}$ Now we randomly select an information structure $\theta_{t}$ which contains the report ${\bm{x}}_{t}$ for every $t$. For simplicity, we also denote ${\bm{x}}={\bm{x}}_{0},{\bm{x}}^{\prime}={\bm{x}}_{T+1},\theta=\theta_{0},\theta^{\prime}=\theta_{T+1}$. By definition,
$S(opt,k)$ $\geq\frac{1}{T+1}\sum_{t=1}^{T+1}|opt_{\theta_{t}}({\bm{x}}_{t})-opt_{\theta_{t-1}}({\bm{x}}_{t-1})|$
$\geq\frac{1}{T+1}|\sum_{t=1}^{T+1}opt_{\theta_{t}}({\bm{x}}_{t})-opt_{\theta_{t-1}}({\bm{x}}_{t-1})|$
$=\frac{1}{T+1}|opt_{\theta^{\prime}}({\bm{x}}^{\prime})-opt_{\theta({\bm{x}})}|$
$=\frac{1}{T+1}S(opt,n)$
Thus $\frac{S(opt,k)}{S(opt,n)}=\frac{1}{T+1}=\Omega(\gamma)$.
Corollary 6.9.
For the linear setting, $R(\Theta,\Sigma)=\Theta(\gamma)$.
7 Conclusion
In this work, we analyze the information aggregation problem under both truthful and adversarial experts. Our results show that for the binary decision problem, the optimal aggregator is a hard sigmoid function under many different settings. In particular, the truncated mean is optimal under mild conditions. We evaluate our aggregators by an ensemble learning task. For the general setting with more flexible information structures and experts’ reports space, we provide some negative results that the optimal aggregator is vulnerable to adversarial experts. In addition, we provide a metric to estimate the regret regarding the benchmark function $opt$.
For future work, it would be interesting to extend the binary decision aggregation problem into a multi-state case, where the adversarial strategies are more diverse. Another possible direction is exploring the performance of the truncated mean in other general information structures, such as the substitute information structure (neyman2022you). Moreover, we wonder what if we connect behavioral economics by considering other types of experts, such as unreliable experts who will make mistakes with probabilities (dawid1979maximum).